# Electroweak Phase Transition in Two Scalar Singlet Model with pNGB Dark Matter

### v1

@@ Abstract
We investigate the dynamics of the electroweak phase transition within an extended Standard Model framework that includes one real scalar $(\Phi)$ and one complex scalar $(S)$, both of which are SM gauge singlets. The global $U(1)$ symmetry is softly broken to a $\mathcal{Z}_{3}$ symmetry by the $S^{3}$ term in the scalar potential. After this $U(1)$ symmetry breaking, the imaginary component of the complex scalar $(S)$ acts as a pseudo-Nambu-Goldstone boson (pNGB) dark matter candidate, naturally stabilized by $\mathcal{Z}_{2}$ symmetry of the scenario. Specially, the spontaneous breaking of the global $U(1)$ symmetry to a discrete $\mathcal{Z}_{3}$ subgroup can introduce effective cubic terms in the scalar potential, which facilitates a strong first-order phase transition. We analyze both single-step and multi-step first-order phase transitions, identifying the parameter space that satisfies the dark matter relic density constraints, complies with all relevant experimental constraints, and exhibits a strong first-order electroweak phase transition. The interplay of these criteria significantly restricts the model parameter space, often leading to an under-abundant relic density. Moreover, we delve into the gravitational wave signatures associated with this framework, offering valuable insights that complement traditional dark matter direct and indirect detection methods.
@@ I Introduction
The detection of gravitational waves by the LIGO collaboration LIGOScientific:2016aoc has steered a new era of precision cosmology, providing a unique window into the physics of the early universe. Gravitational waves, characterized by their distinct sources and detectability Kuroda_2015 , offer invaluable insights into cosmic phenomena. Among these, cosmological first-order phase transitions stand out as one of the promising sources of stochastic gravitational waves via the nucleation of the bubbles of the broken phase. The electroweak phase transition in the Standard Model with the observed Higgs boson mass $\sim 125$ GeV, is characterized as a crossover rather than a first-order phase transition Kajantie:1996mn ; Kajantie:1996qd and thus precludes the generation of the stochastic gravitational waves PhysRevLett.77.2887 . However, various well-motivated extensions to the SM have been proposed that can generate first-order phase transitions and eventually GW signals. Several studies on FOPT and GW has been done in various BSM Models like a single scalar singlet extension Hashino_2017 , 2 scalar singlet extension Shajiee_2019 , 2HDM Goncalves:2021egx . Such GW waves are potentially discoverable by future space-based interferometers, such as LISA LISA:2017pwj ; Baker:2019nia , DECIGO Seto:2001qf ; Kawamura:2006up ; Yagi:2011wg ; Isoyama:2018rjb , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi .
@@
Intriguingly, the interplay between an additional scalar singlet and the Standard Model Higgs doublet can potentially induce two-step phase transitions. However, achieving a strong first-order phase transition (SFOPT) within the confines of a $\mathcal{Z_{\rm 2}}$ symmetric model necessitates a substantial number of additional scalar fields, typically around twelve, as demonstrated in Espinosa_2008 . Furthermore, investigations into phase transitions in conjunction with pseudo-Nambu-Goldstone boson (pNGB) dark matter within models featuring a softly broken global $U(1)$ symmetry by a $\mathcal{Z_{\rm 2}}$ subgroup has revealed that all transitions leading to the true vacuum are of second order Kannike:2019wsn . Consequently, such scenarios are incompatible with the production of gravitational waves.
@@
To induce a first-order phase transition, explicit symmetry-breaking terms, such as soft cubic interactions Kannike:2019mzk or the most general symmetry-breaking operators Alanne:2020jwx , are often introduced at the tree level scalar potential. In particular, the spontaneous breaking of a global $U(1)$ symmetry to a discrete $\mathcal{Z_{\rm 3}}$ subgroup Kannike:2019mzk can generate effective cubic terms in the scalar potential, thereby facilitating a strongly first-order phase transition Witten:1984rs ; Hogan:1983ixn ; Steinhardt:1981ct . In such models, both the broken and unbroken phases stabilize the dark matter candidate due to the symmetry $S\rightarrow S^{\dagger}$ present in the scalar potential where $S$ is the complex singlet scalar. The main difference between the $\mathcal{Z_{\rm 2}}$ and $\mathcal{Z_{\rm 3}}$ is the presence of the cubic terms in the latter which respect the $\mathcal{Z_{\rm 3}}$ symmetry of the potential. Apart from suppressing the direct detection cross section Kannike:2019mzk , the presence of the cubic terms generates a potential barrier at the tree level Ghosh:2022fzp promoting the occurrence of a strong first-order phase transition. The production of gravitational waves (GWs) during such phase transitions has been studied in BSM scenarios with additional scalar fields Espinosa_2012 ; Alanne_2016 ; Cheng_2018 . The phenomenological implications of first-order phase transitions (FOPT) and gravitational wave (GW) production within a $\mathcal{Z_{\rm 3}}$-symmetric model featuring a complex scalar singlet were reported in Kang_2018 ; Kannike:2019mzk .
@@
In our analysis, we first scan the multi-dimensional parameter space that satisfies various theoretical and experimental constraints such as vacuum stability, perturbative unitarity, electroweak precision observables (S & T) parameters, etc. In this framework, we have two real scalars, in addition to the SM Higgs, thus the field configuration space becomes 3-dimensional. This multi-dimensional field configuration makes the shape of the phase transitions more intriguing. We use the publicly available CosmoTransitionsWainwright_2012 package to generate the results of the phase structures of the scalar fields. We find seven benchmark points (BP1-BP7) that show strong first-order phase transitions (SFOPT). Out of these, three benchmark points (BP1, BP2 & BP3) show SFOPT along both the SM and BSM Higgs field directions, while for BP4-BP7, we get SFOPT along the BSM Higgs direction only. As mentioned before, the pNGB dark matter is an integral part of this scenario. Therefore, the next obvious scrutiny would be to check whether the aforementioned seven benchmark points are consistent with the DM relic density, direct and indirect search limits. The seven benchmark points are consistent with dark matter relic density. Most of the points show underabundant relic density except for BP3 and BP7 and thus can be a viable dark matter component. However, it turns out that for some of these benchmark points, the dark matter direct detection cross-section is too large to be satisfied by the current DM direct detection bounds due to the simultaneous requirement of SFOPT and correct relic density which requires high and low values of the scalar quartic couplings respectively. Points where SFOPT is observed and all the latest dark matter experimental results are satisfied, the dark matter relic density is highly underabundant. Further, we show the GW spectrum of those points and whether they would possibly be detected in upcoming future experiments like DECIGO and U-DECIGO.
@@ II The Model
The $\mu_{3}$ term in Eq.(1) softly breaks the global $U(1)$ symmetry. In addition, the potential has discrete symmetry $\mathcal{Z}_{2}$ symmetry, $S\to S^{\dagger}$, just like the original pNGB dark matter model. This $\mathcal{Z}_{2}$ symmetry, as imposed also prevents the odd power terms of $\Phi$ from appearing in the potential.
@@
where, $v_{h},v_{s}$ and $v_{\phi}$ represent vacuum expectation values for the $H,S$ and $\Phi$ fields respectively. Throughout our analysis we assume these vacuum expectation values to be real. The non-zero vev of $S$ field, $\langle S\rangle=v_{s}\neq 0$, spontaneously breaks the $\mathcal{Z}_{3}$ symmetry and the pNGB field $\chi$ transforms as $\chi\rightarrow-\chi$ under the $\mathcal{Z}_{2}$ symmetry and therefore it qualifies as a stable DM candidate even if $\mathcal{Z}_{3}$ is broken.
@@
Note that, $m^{2}_{\chi}$ is proportional to $\mu_{3}$ indicating that the soft breaking of the $U(1)$ symmetry is solely responsible for giving mass to the pNGB dark matter candidate. In the absence of this soft-breaking term, the massive pseudo-Nambu-Goldstone boson (pNGB) would become a massless Nambu-Goldstone boson upon the spontaneous breaking of the global $U(1)$ symmetry.
@@ III Theoretical and Experimental Constraints
$\bullet$ Higgs observables: In the extended Higgs sector, properties of the SM like Higgs should remain intact. In particular, we impose that the lightest mass eigenstate $(H_{1})$ of the three CP-even scalars is identified with the measured SM Higgs with a mass $\approx 125$ GeV and its SM couplings PhysRevD.90.052004 ; Chatrchyan_2014 . In any BSM physics, Higgs signal strengths are used to measure deviations in the production and decay rates of the Higgs boson from the predictions made by the Standard Model (SM). Deviations from the Standard Model (SM) predictions for Higgs boson signal strengths can indicate the presence of new physics phenomena. Conversely, precise measurements of these signal strengths can be used to constrain the parameters of Beyond-the-Standard-Model (BSM) scenarios. In this analysis, we will focus on the latter role of Higgs signal strengths to constrain BSM parameters. In our scenario, the production cross-sections for $H_{1}$ are suppressed by a factor of $(\cos\alpha_{1}\cos\alpha_{2})^{2}$, while the branching ratios remain unchanged. Therefore, one can define the signal strength for each production mode $i$ and decay chain $i\to H_{1}\to f$ as :
@@
where $\sigma_{\rm SM}$ and $\Gamma_{\rm SM}$ denote the SM Higgs boson production and decay rates at the mass $m_{H_{3}}$ and $\psi_{\rm SM}$ represent the SM fermions and gauge bosons. In addition to these decay modes, $H_{3}\to H_{1}H_{1},H_{1}H_{2},~{}~{}{\rm or}~{}~{}H_{2}H_{2}$ are also possible if kinematically allowed. The coupling of $H_{2}$ to the SM particles is doubly suppressed by the factor $\sin\alpha_{1}\cos\alpha_{2}$ and considering the mass hierarchy in this framework, $H_{2}$ can be produced from the decays of $H_{3}$, when it is kinematically accessible. To extract constraints from the SM-like Higgs boson properties we use HiggsSignals 2.2.3 HiggsSignals2 , while to put limits on mixing angles and masses of heavier scalars we use HiggsBounds 5.3.2 HiggsBounds5 . The values of heavy scalar masses $(m_{H_{2}}~{}\&~{}m_{H_{3}})$ and three mixing angles $(\alpha_{1},\alpha_{2},\alpha_{3})$ quoted in our benchmark points satisfy the observed limits at $95\%$ CL.
@@ IV One Loop Finite Temperature Effective Potential
where the partial derivatives are taken with respect to $h$ , $s$ and $\phi$ fields expressed as $h_{a/b}$. The derivatives are evaluated at the vacuum expectation values of the respective fields, $\langle H\rangle=v_{h}$, $\langle S\rangle=v_{s}$, $\langle\Phi\rangle=v_{\phi}$ and $\Delta V$ is the effective potential at zero temperature excluding the tree level part of the potential. The expression of each coefficient of counter terms is shown in Appendix D. The final expression of the total effective potential is given by
@@
The critical temperature $T_{c}$ in a phase transition is determined by setting the potential values at two vacuum expectation values (vevs), namely the High vev and the Low vev, equal to each other. The existence of a barrier between the symmetric phase (high temperature) and the broken phase (low temperature) is the requirement for the strong first order phase transition (SFOPT). This can be established using the following mathematical expression Patel_2011 .
@@
with $\Delta H_{i}$ is the difference of high and low vevs of the SM/BSM scalar field. The criterion for an SFOPT is given by $\zeta_{c,i}\geq 1$. It is important to recognize that the overall effective potential in Eq.(39) explicitly depends on the gauge choice. Consequently, the extrema of the effective potential and the order parameter $\zeta_{c}$ are gauge dependent PhysRevD.9.3320 ; Patel_2011 ; PhysRevD.51.4525 ; Garny_2012 . In our analysis, all finite temperature calculations are performed in the Landau gauge ($\xi=0$) [^1] It is important to note that one can construct gauge independent variables using Nielson’s identities from the effective potential Nielsen:1975fs ; PhysRevD.13.3469 . Detailed analysis is beyond the scope of this paper.. Due to the presence of $\mu_{3}$ term in the tree level potential, the barrier formation takes place in the tree level which facilitates FOPT. Therefore, the gauge dependence that appears through one-loop-induced corrections does not significantly affect our results. Such studies have been conducted in the literature Borah:2023zsb ; Ghosh:2022fzp .
@@
As we shall see later in Sec.VI, the main ingredients to calculate the stochastic gravitational wave spectrum from cosmological phase transitions are the quantities $\{\alpha,\frac{\beta}{H_{n}},T_{n}\}$, where $\alpha$ is related to the latent heat evolved during the phase transition, $\beta$ is the inverse time duration of phase transition, $T_{n}$ is the nucleation temperature defined above in Eq.(45) and $H_{n}$ is the Hubble parameter evaluated at the nucleation temperature. We generate the results of the phase structures of the scalar fields using the publicly available CosmoTransitions Wainwright_2012 package and finally evaluate the output quantities $\{\alpha,\frac{\beta}{H_{n}},T_{n}\}$ corresponding to our benchmark points.

### v2

@@ Abstract
We investigate the dynamics of the electroweak phase transition within an extended Standard Model framework that includes one real scalar $(\Phi)$ and one complex scalar $(S)$, both of which are SM gauge singlets. The global $U(1)$ symmetry is softly broken to a $\mathcal{Z}_{3}$ symmetry by the $S^{3}$ term in the scalar potential. After this $U(1)$ symmetry breaking, the imaginary component of the complex scalar $(S)$ acts as a pseudo-Nambu-Goldstone boson (pNGB) dark matter candidate, naturally stabilized by the $\mathcal{Z}_{2}$ symmetry of the scenario. Specially, the spontaneous breaking of the global $U(1)$ symmetry to a discrete $\mathcal{Z}_{3}$ subgroup can introduce effective cubic terms in the scalar potential, which facilitates a strong first-order phase transition. We analyze both single-step and multi-step first-order phase transitions, identifying the parameter space that satisfies the dark matter relic density constraints, complies with all relevant experimental constraints, and exhibits a strong first-order electroweak phase transition. The interplay of these criteria significantly restricts the model parameter space, often leading to an underabundant relic density. Moreover, we delve into the gravitational wave signatures associated with this framework, offering valuable insights that complement traditional dark matter direct and indirect detection methods.
@@ I Introduction
The detection of gravitational waves by the LIGO collaboration LIGOScientific:2016aoc has steered a new era of precision cosmology, providing a unique window into the physics of the early universe. Gravitational waves, characterized by their distinct sources and detectability Kuroda:2015owv , offer invaluable insights into cosmic phenomena. Among these, cosmological first-order phase transitions stand out as one of the promising sources of stochastic gravitational waves via the nucleation of the bubbles of the broken phase. The electroweak phase transition in the Standard Model with the observed Higgs boson mass $\sim 125$ GeV, is characterized as a crossover rather than a first-order phase transition PhysRevLett.77.2887 ; Kajantie:1996qd and thus precludes the generation of the stochastic gravitational waves PhysRevLett.77.2887 . However, various well-motivated extensions to the SM have been proposed that can generate first-order phase transitions and eventually GW signals. Several studies on FOPT and GW has been done in various BSM Models like a single scalar singlet extension Hashino:2016xoj , 2 scalar singlet extension Shajiee:2018jdq ; Ghorbani:2019itr ; Ghorbani:2024twk , 2HDM Goncalves:2021egx . Such GW waves are potentially discoverable by future space-based interferometers, such as LISA LISA:2017pwj ; Baker:2019nia , DECIGO Seto:2001qf ; Kawamura:2006up ; Yagi:2011wg ; Isoyama:2018rjb , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi .
@@
Intriguingly, the interplay between an additional scalar singlet and the Standard Model Higgs doublet can potentially induce two-step phase transitions. However, achieving a strong first-order phase transition (SFOPT) within the confines of a $\mathcal{Z_{\rm 2}}$ symmetric model necessitates a substantial number of additional scalar fields, typically around twelve, as demonstrated in PhysRevD.78.123528 . Furthermore, investigations into phase transitions in conjunction with pNGB dark matter within models featuring a softly broken global $U(1)$ symmetry by a $\mathcal{Z_{\rm 2}}$ subgroup has revealed that all transitions leading to the true vacuum are of second order Kannike:2019wsn . Consequently, such scenarios are incompatible with the production of gravitational waves.
@@
To induce a first-order phase transition, explicit symmetry-breaking terms, such as soft cubic interactions Kannike:2019mzk or the most general symmetry-breaking operators Alanne:2020jwx , are often introduced at the tree level scalar potential. In particular, the spontaneous breaking of a global $U(1)$ symmetry to a discrete $\mathcal{Z_{\rm 3}}$ subgroup Kannike:2019mzk can generate effective cubic terms in the scalar potential, thereby facilitating a strongly first-order phase transition Witten:1984rs ; Hogan:1983ixn ; Steinhardt:1981ct . In such models, both the broken and unbroken phases stabilize the dark matter candidate due to the symmetry $S\rightarrow S^{\dagger}$ present in the scalar potential where $S$ is the complex singlet scalar. The main difference between the $\mathcal{Z_{\rm 2}}$ and $\mathcal{Z_{\rm 3}}$ is the presence of the cubic terms in the latter which respect the $\mathcal{Z_{\rm 3}}$ symmetry of the potential. Apart from suppressing the direct detection cross section Kannike:2019mzk , the presence of the cubic terms generates a potential barrier at the tree level Ghosh:2022fzp promoting the occurrence of a strong first-order phase transition. The production of gravitational waves (GWs) during such phase transitions has been studied in BSM scenarios with additional scalar fields Espinosa:2011ax ; Alanne:2016wtx ; Cheng:2018ajh . The phenomenological implications of first-order phase transitions and gravitational wave production within a $\mathcal{Z_{\rm 3}}$-symmetric model featuring a complex scalar singlet were reported in Kang:2017mkl ; Kannike:2019mzk ; Chiang:2019oms .
@@
In our analysis, we first scan the multi-dimensional parameter space that satisfies various theoretical and experimental constraints such as vacuum stability, perturbative unitarity, electroweak precision observables (S & T) parameters, etc. In this framework, we have two real scalars, in addition to the SM Higgs, thus the field configuration space becomes 3-dimensional. This multi-dimensional field configuration makes the shape of the phase transitions more intriguing. We use the publicly available CosmoTransitionsWainwright:2011kj package to generate the results of the phase structures of the scalar fields. We find seven benchmark points (BP1-BP7) that show strong first-order phase transitions. Out of these, three benchmark points (BP1, BP2 & BP3) show SFOPT along both the SM and BSM Higgs field directions, while for BP4-BP7, we get SFOPT along the BSM Higgs direction only. As mentioned before, the pNGB dark matter is an integral part of this scenario. Therefore, the next obvious scrutiny would be to check whether the aforementioned seven benchmark points are consistent with the DM relic density, direct and indirect search limits. The seven benchmark points are consistent with dark matter relic density. Most of the points show underabundant relic density except for BP3 and BP7 and thus can be a viable dark matter component. However, it turns out that for some of these benchmark points, the dark matter direct detection cross-section is too large to be satisfied by the current DM direct detection bounds due to the simultaneous requirement of SFOPT and correct relic density which requires high and low values of the scalar quartic couplings respectively. Points where SFOPT is observed and all the latest dark matter experimental results are satisfied, the dark matter relic density is highly underabundant. Further, we show the GW spectrum of those points and whether they would possibly be detected in upcoming future experiments like DECIGO and U-DECIGO.
@@ II The Model
The $\mu_{3}$ term in Eq.(1) softly breaks the global $U(1)$ symmetry. In addition, the potential has discrete symmetry $\mathcal{Z}_{2}$ symmetry, $S\to S^{\dagger}$, just like the original pNGB dark matter model. This $\mathcal{Z}_{2}$ symmetry, as imposed, also prevents the odd power terms of $\Phi$ from appearing in the potential.
@@
where, $v_{h},v_{s}$ and $v_{\phi}$ represent vacuum expectation values for the $H,S$ and $\Phi$ fields respectively. Throughout our analysis we assume these vacuum expectation values to be real. The non-zero vev of $S$ field, $\langle S\rangle=v_{s}\neq 0$, spontaneously breaks the $\mathcal{Z}_{3}$ symmetry and the pNGB field $\chi$ transforms as $\chi\rightarrow-\chi$ under the $\mathcal{Z}_{2}$ symmetry and therefore it qualifies as a stable DM candidate even if $\mathcal{Z}_{3}$ is broken.  [^1] In general, spontaneous breaking of discrete $\mathcal{Z_{\rm 3}}$ symmetry leads to the formation of domain walls. However, this the issue of domain wall creation can be avoided if $U(1)$ symmetry is gauged in a UV complete model Gross:2017dan ; Abe:2022mlc ; Dvali:1994wv .
@@
Note that, $m^{2}_{\chi}$ is proportional to $\mu_{3}$ indicating that the soft breaking of the $U(1)$ symmetry is solely responsible for giving mass to the pNGB dark matter candidate. In the absence of this soft-breaking term, the massive pNGB would become a massless Nambu-Goldstone boson upon the spontaneous breaking of the global $U(1)$ symmetry.
@@ III Theoretical and Experimental Constraints
$\bullet$ Higgs observables: In the extended Higgs sector, properties of the SM like Higgs should remain intact. In particular, we impose that the lightest mass eigenstate $(H_{1})$ of the three CP-even scalars is identified with the measured SM Higgs with a mass $\approx 125$ GeV and its SM couplings PhysRevD.90.052004 ; CMS:2013fjq . In any BSM physics, Higgs signal strengths are used to measure deviations in the production and decay rates of the Higgs boson from the predictions made by the SM. Deviations from the SM predictions for Higgs boson signal strengths can indicate the presence of new physics phenomena. Conversely, precise measurements of these signal strengths can be used to constrain the parameters of BSM scenarios. In this analysis, we will focus on the latter role of Higgs signal strengths to constrain BSM parameters. In our scenario, the production cross-sections for $H_{1}$ are suppressed by a factor of $(\cos\alpha_{1}\cos\alpha_{2})^{2}$, while the branching ratios remain unchanged. Therefore, one can define the signal strength for each production mode $i$ and decay chain $i\to H_{1}\to f$ as :
@@
where $\sigma_{\rm SM}$ and $\Gamma_{\rm SM}$ denote the SM Higgs boson production and decay rates at the mass $m_{H_{3}}$ and $\psi_{\rm SM}$ represent the SM fermions and gauge bosons. In addition to these decay modes, $H_{3}\to H_{1}H_{1},H_{1}H_{2},~{}~{}{\rm or}~{}~{}H_{2}H_{2}$ are also possible if kinematically allowed. The coupling of $H_{2}$ to the SM particles is doubly suppressed by the factor $\sin\alpha_{1}\cos\alpha_{2}$ and considering the mass hierarchy in this framework, $H_{2}$ can be produced from the decays of $H_{3}$, when it is kinematically accessible. To extract constraints from the SM-like Higgs boson properties we use HiggsSignals 2.2.3 HiggsSignals2 , while to put limits on mixing angles and masses of heavier scalars we use HiggsBounds 5.3.2 Bechtle:2020pkv . The values of heavy scalar masses $(m_{H_{2}}~{}\&~{}m_{H_{3}})$ and three mixing angles $(\alpha_{1},\alpha_{2},\alpha_{3})$ quoted in our benchmark points satisfy the observed limits at $95\%$ CL.
@@ IV One Loop Finite Temperature Effective Potential
where the partial derivatives are taken with respect to $h$ , $s$ and $\phi$ fields expressed as $h_{a/b}$. The derivatives are evaluated at the vacuum expectation values of the respective fields, $\langle H\rangle=v_{h}$, $\langle S\rangle=v_{s}$, $\langle\Phi\rangle=v_{\phi}$, and $\Delta V$ is the effective potential at zero temperature excluding the tree-level part of the potential. The expression of each coefficient of counter terms is shown in Appendix D. The final expression of the total effective potential is given by
@@
The critical temperature $T_{c}$ in a phase transition is determined by setting the potential values at two vacuum expectation values (vevs), namely the High vev and the Low vev, equal to each other. The existence of a barrier between the symmetric phase (high temperature) and the broken phase (low temperature) is the requirement for the strong first order phase transition (SFOPT). This can be established using the following mathematical expression Patel:2011th .
@@
with $\Delta H_{i}$ is the difference of high and low vevs of the SM/BSM scalar field. The criterion for an SFOPT is given by $\zeta_{c,i}\geq 1$. It is important to recognize that the overall effective potential in Eq.(39) explicitly depends on the gauge choice. Consequently, the extrema of the effective potential and the order parameter $\zeta_{c}$ are gauge dependent PhysRevD.9.3320 ; Patel:2011th ; PhysRevD.51.4525 ; Garny:2012cg . In our analysis, all finite temperature calculations are performed in the Landau gauge ($\xi=0$) [^2] It is important to note that one can construct gauge independent variables using Nielson’s identities from the effective potential Nielsen:1975fs ; PhysRevD.13.3469 . Detailed analysis is beyond the scope of this paper.. Due to the presence of the $\mu_{3}$ term in the tree-level potential, barrier formation takes place at the tree level which facilitates FOPT. Therefore, the gauge dependence that appears through one-loop-induced corrections does not significantly affect our results. Such studies have been conducted in the literature Borah:2023zsb ; Ghosh:2022fzp .
@@
As we shall see later in Sec.VI, the main ingredients to calculate the stochastic gravitational wave spectrum from cosmological phase transitions are the quantities $\{\alpha,\frac{\beta}{H_{n}},T_{n}\}$, where $\alpha$ is related to the latent heat evolved during the phase transition, $\beta$ is the inverse time duration of phase transition, $T_{n}$ is the nucleation temperature defined above in Eq.(45) and $H_{n}$ is the Hubble parameter evaluated at the nucleation temperature. We generate the results of the phase structures of the scalar fields using the publicly available CosmoTransitions Wainwright:2011kj package and finally evaluate the output quantities $\{\alpha,\frac{\beta}{H_{n}},T_{n}\}$ corresponding to our benchmark points.